# Rigidity transition in polydisperse shear-thickening suspensions

### v1

@@ II Modeling and simulations
All the simulations, if not mentioned otherwise, are carried out with $N=1000$ spherical particles, interacting with static friction coefficient $\mu=0.5$ at contact. We have studied the systems in the limit of high non-dimensional stress ($\sigma=\hat{\sigma}/\sigma_{0}$) values of 25 and 100, where shear jamming is nearly independent of the applied stress [50, 55]. For consistency, all the length scales in the simulations are non-dimensionalized by a characteristic particle radius $a=1.29$. To simulate polydisperse suspensions we have considered normal and log-normal particle size distributions. The mean ($\langle a\rangle$), standard deviation ($s.d.$) and skewness ($s$) of the distributions are defined as, $\langle a\rangle=\displaystyle\frac{1}{N}\sum\limits_{i}a_{i}$, $s.d.=\left(\displaystyle\frac{1}{N}\sum\limits_{i}a_{i}^{2}-\langle a\rangle^{2}\right)^{1/2}$, and $s=\left(\displaystyle\frac{1}{N}\sum\limits_{i}(a_{i}-\langle a\rangle)^{3}\right)/s.d.^{3}$, respectively. Typically, we refer to polydisperse system by its polydispersity index $\alpha$, defined as $\alpha=s.d./\langle a\rangle$. The details of the particle size distributions simulated in the present study is tabulated in Table 1.
@@
Figure 2: Divergence of relative viscosity $\eta_{r}$ as a function of packing fraction $\phi$ for (a) normal and its statistically equivalent bidisperse distributions with polydispersity index $\alpha=0.133$, at stress $\sigma=25$ and for (b) log-normal and its statistically equivalent bidisperse distributions, with $\alpha=0.374$, at $\sigma=100$. The lines represent M-P fit: $\eta_{r}\sim\left(1-\frac{\phi}{\phi_{J}(\sigma)}\right)^{-2}$, where $\phi_{J}(\sigma)$ is the packing fraction in the shear jamming limit.
III Rigidity analysis and validation study
The identification of the rigid clusters from a given simulation snapshot is performed by applying the (3,3) pebble game (PG) algorithm, which is a constraint counting algorithm to identify the rigid nodes from a given graph [25, 37]. Here, each particle is assigned with 3 pebbles corresponding to 2 translational and 1 rotational degrees of freedom. In a contact network, whenever a pair of particles are in frictional contact, one of the particle loses one or two pebbles, depending on whether the contact is sliding or non-sliding, respectively, resulting in addition of constraints. Likewise, for a given cluster of particles, all such constraints are constructed and the number of free pebbles are counted. If only a total of three free pebbles remain for the cluster, then the cluster is designated as a rigid one according to PG algorithm. Identification and growth of such rigid clusters have been considered in characterizing jamming in dry granular media [24, 15, 2] and also in understanding the mechanism of shear jamming in dense suspensions [59, 51]. In this work we have investigated the effects of polydispersity on the critical transition of cluster rigidity in terms of an order-parameter defined by the fraction of particles in rigid cluster, $f_{\text{rig}}$, same as that considered by Santra et al. [51] in there study on bidisperse systems.
$f_{\text{rig}}=\langle\frac{1}{N}\sum\limits_{i=1}^{N}n_{i}\rangle$ (4)
$\chi_{\text{rig}}=N\langle(\frac{1}{N}\sum\limits_{i=1}^{N}n_{i}-f_{\text{rig}})^{2}\rangle$ (5)
In the subsequent section we discuss the scaling behaviour of $f_{\text{rig}}$ and $\chi_{\text{rig}}$ for the polydisperse systems and compared that with the bidisperse cases.
To validate the bulk rheological behaviour of a 2D polydisperse suspension, we have compared their viscosity with that of a statistically equivalent bidisperse suspension. Here, two suspensions are defined to be statistically equivalent if $\langle a\rangle$, $s.d.$ and $s$ are identical for both the systems, as illustrated in Figs. 1 (a) and 1 (b). In Fig. 2 we show that the divergence of the relative viscosity ($\eta_{r}=\eta_{\text{suspension}}/\eta_{0}$) as a function of the packing fraction is identical for the polydisperse and statistically equivalent bidisperse suspensions in 2D at stresses $\sigma=25$ and $\sigma=100$, for both normal and log-normal distributions. This is in agreement with that observed in case of the suspensions simulated in 3D [43]. Furthermore, Figs. 2 (a) and 2 (b) demonstrate that the viscosity divergence can be described by the Maron-Pierce (M-P) model ($\eta_{r}\sim\left(1-\frac{\phi}{\phi_{J}(\sigma)}\right)^{-2}$, where $\phi_{J}(\sigma)$ is the shear jamming fraction), similar to a 3D suspension [32, 43].
@@ IV.1 Scaling at the rigidity transition
Figure 3: Variation of the order-parameter $f_{\text{rig}}$ with solid packing fraction $\phi$ for normal and its statistically equivalent bidisperse distributions with $\alpha=0.133$, at stress (a) $\sigma=25$ and (b) $\sigma=100$. (c) Variation of $f_{\text{rig}}$ with $\phi$ for log-normal and its statistically equivalent bidisperse distributions with $\alpha=0.374$, at $\sigma=100$. $\phi_{c}$ denotes the corresponding critical packing fraction. The solid lines indicate the fit analogous to the 2D-Ising model with scaling exponent $\beta=1/8$.
@@
Figure 4: Variation of the susceptibility $\chi_{\text{rig}}$ with solid packing fraction $\phi$ for normal and its statistically equivalent bidisperse distributions with $\alpha=0.133$, at stress (a) $\sigma=25$ and (b) $\sigma=100$. (c) Variation of $\chi_{\text{rig}}$ with $\phi$ for log-normal and its statistically equivalent bidisperse distributions with $\alpha=0.374$, at $\sigma=100$. The vertical line at $\phi=\phi_{c}$ indicates the critical point at which $\chi_{\text{rig}}$ diverges.
In this section, we characterize the growth of rigid clusters near shear jamming by investigating the scaling of the order-parameter $f_{\text{rig}}$ and susceptibility $\chi_{\text{rig}}$. Similar to bidisperse suspensions [51], Figs. 3 show that $f_{\text{rig}}$ for polydisperse systems also increases with solid packing fraction $\phi$ and undergoes a sharp transition at $\phi=\phi_{c}$, where $\phi_{c}(<\phi_{J})$ denotes the critical packing fraction. Interestingly, the variations of $f_{\text{rig}}$ with $\phi$ for the polydisperse suspensions are exactly identical to the statistically equivalent bidisperse systems. Furthermore, the growth of $f_{\text{rig}}$ above $\phi_{c}$ can be well fitted by the scaling relation, $f_{\text{rig}}\sim[\phi-\phi_{c}]^{\beta}$, analogous to that for the critical transition in 2D-Ising model, with $\beta=1/8$ as the scaling exponent [26, 63, 31]. Here, $f_{\text{rig}}$ and $\phi$ are the variables analogous to magnetization and temperature, respectively, in critical transition of a magnetic system [26]. The critical transition is further demonstrated in Figs. 4 by the presence of a sharp peak in the variation of $\chi_{\text{rig}}$ as a function of $\phi$, for both polydisperse and bidisperse systems. Similar to the order-parameter, the susceptibility variation is also identical for the polydisperse and the statistically equivalent bidisperse suspensions. Such identical variations of the order-parameter and susceptibility clearly indicates that not only the bulk rheology but also the growth and development of the flow microstructure in a shear thickening suspension of polydisperse particles can be effectively predicted from the microstructure of a relatively simple, equivalent bidisperse system. This implies that the microscopic stress environment surrounding a particle in a polydisperse system is same as that in the corresponding statistically equivalent bidisperse suspension, reasserting the claim by Pednekar et al. [43]. The similarity between the microstructural features of a polydisperse suspension with that of its statistically equivalent bidisperse system is further demonstrated in the Appendix by the identical behaviour of gradient direction velocity correlation functions.
Moreover, we can construct a master-curve, as shown in Fig. 5 (a), to collapse the values of the order-parameter over a range of packing fraction, at different conditions of polydispersity, by plotting $f_{\text{rig}}$ as a function of the relative distance from the critical point ($\phi-\phi_{c}$). The susceptibility data, as presented in Fig. 5 (b), are also found to collapse on a master-curve when plotted as a function of $|\phi-\phi_{c}|$. Interestingly, close to the critical point, i.e. $|\phi-\phi_{c}|<10^{-2}$, the susceptibility follows the scaling, $\chi_{\text{rig}}\sim|\phi-\phi_{c}|^{-\gamma}$, with $\gamma=7/4$, which corresponds to the scaling exponent for susceptibility in 2D-Ising model [63, 31]. We have further verified this scaling exponent from finite size scaling analysis discussed in the following section.
@@
Figure 5: Variation of (a) the order-parameter $f_{\text{rig}}$ and (b) susceptibility $\chi_{\text{rig}}$ as a function of the relative distance from the critical point ($\phi-\phi_{c}$) for different conditions of polydispersity at stress values of 100 and 25.
@@ IV.2 Finite size scaling
Figure 6: Pair-correlation function for rigid particles at (a) $\phi=0.78$ and (b) $0.794$ for normally distributed polydisperse systems ($\alpha=0.133$) with number of particles $N=1000$, $2000$ and $3000$. All the data are computed at stress $\sigma=100$.
Critical phase transitions in equilibrium systems are always associated with a diverging correlation length [47, 62, 9, 6] which results into the universal scaling relations for the order-parameter and susceptibility as discussed in the previous section. While our systems of shear thickening suspensions are fundamentally out of equilibrium, the observation of a critical transition for such a system is in itself an interesting finding. Furthermore, the agreement with the scaling exponents of the 2D-Ising model suggests the applicability of such model for explaining even certain non-equilibrium phase transition, such as, the shear jamming at high stress as in the present study, where the system undergoes a transition from a flowable to a non-flowable state. While approaching a critical point, the correlation length ($\xi$), associated with the order parameter, is found to increase as $\xi\sim|t|^{-\nu}$, where $\nu$ is a critical exponent and $|t|$ is defined as the relative distance from the critical point in the limit of an infinite system size, i.e. in the context of the present study $|t|=|\phi-\phi_{c_{\infty}}|$, where $\phi_{c_{\infty}}$ is the critical packing fraction for an infinitely large system [47, 9]. This implies that for a system with linear dimension $L$, the correlation length at the critical transition becomes $\xi=L$, as $L\rightarrow\infty$. For finite size systems, the divergence of the correlation length near the critical point is characterized by the following finite size scaling relations [47, 62, 9, 4, 19, 60],
$\xi(t,L)$ $=$ $L\mathcal{G}(t\,L^{1/\nu})$ (6)
$\chi(t,L)$ $=$ $L^{\gamma/\nu}\mathcal{F}(t\,L^{1/\nu})$ (7)
Here, $\mathcal{G}$ and $\mathcal{F}$ are the scaling functions for the correlation length and susceptibility, respectively, and $t\,L^{1/\nu}$ is the corresponding scaling variable. This allows us to determine the critical exponents, $\gamma$ and $\nu$, from simulations of systems with finite domain and to identify the characteristics of the critical phenomena. Here, we have investigated finite size scaling behaviour of polydisperse suspensions near the rigidity transition by simulating systems with different number of particles ($N=1000$, $2000$ and $3000$) at stress $\sigma=100$. In our 2D simulations, since $N\propto L^{2}$, the system size is defined by $\sqrt{N}$. The correlation length for the rigid clusters can be estimated from a length scale associated with the cluster size. We have evaluated a length scale from the decay of the pair-correlation function, $g(r)$, for the rigid particles (identified from the pebble game algorithm) and used it as an estimate of the correlation length $\xi$ at different system sizes. This is illustrated in Figs. 6, where the variations of $g(r)$ are shown for $\phi=0.78$ and $\phi=0.794$ (where $\phi_{c}\approx 0.794$), at different system sizes. It is to be noted that for any given value of $\phi$, the pair-correlation function is normalized by the corresponding number density of rigid particles only (i.e. non-rigid particles are excluded from the calculation). The long-range correlation length scale is extracted from the decay of the pair-correlation function by identifying the value of $r$ (radial distance) at which $g(r)$ settles to 1, which corresponds to the average number density of rigid particles, as shown in Figs. 6. Interestingly, the correlation length ($\xi$) extracted in this manner is found to increase with increasing system size and packing fraction. This suggests that there is a divergence in the correlation length as the packing fraction approaches the critical point ($\phi_{c}$) from below and at $\phi=\phi_{c}$, the correlation length $\xi\rightarrow\infty$ in the limit of infinite system size.
@@
The finite size scaling relations in Eqs. (3) and (4) are tested for the normally-distributed polydisperse suspensions (with $\alpha=0.133$) by computing $\xi$ and $\chi_{\text{rig}}$ for systems with $N=1000$, $N=2000$ and $N=3000$ particles. Fig. 7 (a) presents the scaling of $\xi/L$ as a function of the scaling variable $t\,L^{1/\nu}$, where $t=(\phi-\phi_{c_{\infty}})$ and $\phi_{c_{\infty}}\approx 0.794$. A good data collapse is observed for $\xi/L$ with a value of the scaling exponent $\nu=1.64$. As shown in Fig. 7 (b), the scaling of $\chi_{\text{rig}}/L^{\gamma/\nu}$ also indicates a good collapse of data with different system sizes. The corresponding values of the scaling exponents are found to be $\gamma=7/4$ and $\nu=1.64$, which are consistent with the susceptibility scaling exponent shown in Fig. 5 (b) and finite size scaling of correlation length in Fig. 7 (a). Interestingly, while the values of the critical exponents $\beta$ and $\gamma$ are consistent with that of the 2D-Ising model, exponent $\nu$ deviates from the value $\nu=1$ obtained in 2D-Ising model. Nevertheless, $\nu=1.64$ is in agreement with the value of the critical exponent in 2D percolation transition for continuum off-lattice systems [57, 8, 42], suggesting that the rigidity transition in shear thickening suspensions can be correlated to a percolation transition [59, 21].
Refer to caption Refer to caption (a) (b)

### v2

@@ III Rigidity analysis and validation study
To validate the bulk rheological behavior of a 2D polydisperse suspension, we have compared their viscosity with that of a statistically equivalent bidisperse suspension. Here, two suspensions are defined to be statistically equivalent if $\langle a\rangle$, $s.d.$ and $s$ are identical for both the systems, as illustrated in Figs. 2 (a) and 2 (b). It is worth noting here that we have considered a truncated normal and log-normal distribution to keep the size of the smallest particle well above $1\,\mu m$, so that the Brownian forces are still negligible. In Fig. 3 we show that the divergence of the relative viscosity ($\eta_{r}=\eta_{\text{suspension}}/\eta_{0}$) as a function of the packing fraction is identical for the polydisperse and statistically equivalent bidisperse suspensions in 2D at stresses $\sigma=25$ and $\sigma=100$, for both normal and log-normal distributions. This is in agreement with that observed in case of the suspensions simulated in 3D [48]. Furthermore, Fig. 3 (a) and 3 (b) demonstrate that the viscosity divergence can be described by the Maron-Pierce (M-P) model ($\eta_{r}\sim\left(1-\frac{\phi}{\phi_{J}(\sigma)}\right)^{-b}$, where $\phi_{J}(\sigma)$ is the shear jamming fraction and $b$ is the scaling exponent obtained by fitting the simulation data), similar to a 3D suspension [38, 48]. The values of $\phi_{J}$ and $b$ estimated for the data in Fig. 3 (a) and Fig. 3 (b) are $\phi_{J}=0.8221\pm 0.0017$, $b=2.0\pm 0.01$, and $\phi_{J}=0.8231\pm 0.0012$, $b=2.03\pm 0.04$, respectively. It is to be noted that while the M-P model fit the whole set of data for the normal distribution, in case of log-normal distribution the data far from the jamming point (i.e. $\phi\ll\phi_{J}$) deviate from the M-P line, as shown in the insets of Fig. 3 (a) and (b).
@@
Figure 3: Divergence of relative viscosity $\eta_{r}$ as a function of packing fraction $\phi$ for (a) normal and its statistically equivalent bidisperse distributions with polydispersity index $\alpha=0.133$, at stress $\sigma=25$ and for (b) log-normal and its statistically equivalent bidisperse distributions, with $\alpha=0.374$, at $\sigma=100$. The lines represent M-P fit: $\eta_{r}\sim\left(1-\frac{\phi}{\phi_{J}(\sigma)}\right)^{-b}$, where $\phi_{J}(\sigma)$ is the packing fraction in the shear jamming limit and $b$ is the scaling exponent. The estimated values of $\phi_{J}$ in sub-figure (a) and (b) are $0.8221\pm 0.0017$ and $0.8231\pm 0.0012$, respectively. The M-P fit is illustrated in the inset by plotting $\eta_{r}$ as a function of $(1-\phi/\phi_{J})$ on a log-log scale.
@@ IV.1 Scaling at the rigidity transition
Figure 4: Variation of the order-parameter $f_{\text{rig}}$ with solid packing fraction $\phi$ for normal and its statistically equivalent bidisperse distributions with $\alpha=0.133$, at stress (a) $\sigma=25$ and (b) $\sigma=100$. (c) Variation of $f_{\text{rig}}$ with $\phi$ for log-normal and its statistically equivalent bidisperse distributions with $\alpha=0.374$, at $\sigma=100$. $\phi_{c}$ denotes the corresponding critical packing fraction. The insets show $f_{\text{rig}}$ as a function of $(\phi-\phi_{c})$ on a log-log scale with the solid line representing a power-law fit with exponent $\beta=0.164\pm 0.025$.
@@
Figure 5: Variation of the susceptibility $\chi_{\text{rig}}$ with solid packing fraction $\phi$ for normal and its statistically equivalent bidisperse distributions with $\alpha=0.133$, at stress (a) $\sigma=25$ and (b) $\sigma=100$. (c) Variation of $\chi_{\text{rig}}$ with $\phi$ for log-normal and its statistically equivalent bidisperse distributions with $\alpha=0.374$, at $\sigma=100$. The vertical line at $\phi=\phi_{c}$ indicates the critical point at which $\chi_{\text{rig}}$ maximizes. The insets represent the corresponding data in linear scale.
In this section, we characterize the growth of rigid clusters near shear jamming by investigating the scaling of the order-parameter $f_{\text{rig}}$ and susceptibility $\chi_{\text{rig}}$. Similar to bidisperse suspensions [56], Figs. 4 show that $f_{\text{rig}}$ for polydisperse systems also increases with solid packing fraction $\phi$ and undergoes a sharp transition at $\phi=\phi_{c}$, where $\phi_{c}(<\phi_{J})$ denotes the critical packing fraction. Notably, the variations of $f_{\text{rig}}$ with $\phi$ for the polydisperse suspensions are identical to the statistically equivalent bidisperse systems. Furthermore, the growth of $f_{\text{rig}}$ above $\phi_{c}$ can be well fitted by the scaling relation, $f_{\text{rig}}\sim[\phi-\phi_{c}]^{\beta}$, analogous to a critical transition where $f_{\text{rig}}$ is the order parameter, $\phi$ is the control parameter and and $\phi_{c}$ denotes the critical point [64, 63, 5, 30, 71, 36]. The critical transition is further demonstrated in Fig. 5 by the presence of a sharp peak in the variation of $\chi_{\text{rig}}$ as a function of $\phi$, for both polydisperse and bidisperse systems. The critical volume fraction $\phi_{c}$ is estimated by fitting a second order polynomial $g(x)$ to the $\log(\chi_{\text{rig}})$ versus $\phi$ data close to the peak, as indicated in Fig. 5, and subsequently equating $g^{\prime}(x)=0$ to evaluate the value of $x(=\phi_{c})$ at which $\chi_{\text{rig}}$ becomes maximum. The uncertainty in the values of $\phi_{c}$ is determined from the estimated error in the fitting parameters and is found to be negligible (order of $\sim 10^{-5}$). The exponent $\beta$ is evaluated by a linear fit to the $f_{\text{rig}}$ data on a log-log scale as indicated in the insets of Fig. 4. It is to be noted that we have evaluated $\beta$ by fitting $f_{\text{rig}}$ for all the different cases of polydisperse suspensions, taken together (this is further discussed in the context of Fig. 7 (a), in the subsequent paragraphs). Notably, the estimated value of $\beta$ is $0.164\pm 0.025$, which is in agreement with the value of critical exponent $\beta=5/36\approx 0.14$ for standard 2D percolation [63, 64, 5]. Recently, Goyal et al. [22] studied the percolation of 4-neighbor particles at DST in 3D systems and found the percolation probability (order parameter) to grow with an exponent of $\beta\approx 0.18$, which is far from the theoretical prediction [63, 29]. On the contrary, considering $f_{\text{rig}}$ as the order parameter for a 2D system we show that there is a critical transition near shear jamming which could be well described by 2D percolation theory. Similar to the order-parameter, the susceptibility variation is also identical for the polydisperse and the statistically equivalent bidisperse suspensions. Such identical variations of the order-parameter and susceptibility clearly indicate that not only the bulk rheology but also the growth and development of the flow microstructure in a shear thickening suspension of polydisperse particles can be effectively predicted from the microstructure of a relatively simple, equivalent bidisperse system. Equivalence in the maximum packing fraction and macroscopic rheological properties between different forms of polydisperse suspensions, mostly in the limit of frictionless jamming are already well explored [15, 69]. However, such equivalence in the microstructural properties, such as cluster distribution or rigidity transitions were not so well reported. Previously, Pednekar,Sidhant et al. [48] have shown that particles of a given size identically contributes to the hydrodynamic viscosity for two statistically equivalent systems, suggesting a similar stress environment for a particle of a given size in polydisperse and equivalent bidisperse suspensions. However, its consequence on the cluster rigidity and critical transition near shear jamming was not well understood. In the present study we have investigated these aspects in the framework of 2D percolation transition. Additionally, the simulation results suggest that at a coarse grained level the local stress becomes insensitive to the local particle composition. This is illustrated in Fig. 6, where we have presented simulation snapshots of a bidisperse and statistically equivalent polydisperse system to show the comparison of the translational velocity distribution (which is related to the local stress distribution) in their respective rigid zones. As shown in Fig. 6 (a) and 6 (b), for both the systems, magnitude of the velocity varies within the same range. Notably, at similar locations (highlighted by blue, green and magenta circles) within the simulation box, the velocity distributions of the particles for both the systems are nearly identical even though the local particle composition in the highlighted regions are very different. It is to be noted that since Fig. 6 displays a single simulation snapshot, exact match in the velocity distribution between the statistically equivalent systems is not observed, for which ensemble averaging over large number of snapshots is required. The similarity between the microstructural features of a polydisperse suspension with that of its statistically equivalent bidisperse system is further demonstrated in the Appendix by the identical behavior of gradient direction velocity correlation functions.
@@
Figure 6: Snapshots of translational velocity distribution of the particles in rigid clusters shown by a colour map based on the magnitude of the particle velocity for (a) bidisperse suspension with $\alpha=0.133$, and (b) statistically equivalent polydisperse suspensions, at $\sigma=100$, $N=1000$. Blue, green and magenta circles are drawn to highlight similar locations in the simulation box for the two equivalent systems.
Refer to caption Refer to caption (a) (b)
Figure 7: Variation of (a) the order-parameter $f_{\text{rig}}$ and (b) susceptibility $\chi_{\text{rig}}$ as a function of the relative distance from the critical point ($\phi-\phi_{c}$) for different conditions of polydispersity at stress values of 100 and 25, for $N=1000$. (a) The inset presents $f_{\text{rig}}$ data on a log-log scale. The solid line represents a fit to the scaling relation $f_{\text{rig}}\sim(\phi-\phi_{c})^{\beta}$, where $\beta=0.164\pm 0.025$. (b) The inset presents $\chi_{\text{rig}}$ data for $N=8000$ particles. The dashed line represents a power law with exponent $3.75$ for both the main plot and inset. The solid line in the main plot and inset represents a power law with exponent $\gamma\approx 0.3$ and $\gamma\approx 2.4$, respectively.
Moreover, we can construct a master-curve, as shown in Fig. 7 (a), to collapse the values of the order-parameter over a range of packing fraction, at different conditions of polydispersity, by plotting $f_{\text{rig}}$ as a function of the relative distance from the critical point ($\phi-\phi_{c}$). The inset in Fig. 7 (a) further demonstrates the data collapse on a log-log scale. As mentioned in the previous paragraph, the data set is fitted on a log-log scale to extract a value of the power law exponent $\beta$, which is found to be $0.164\pm 0.025$, consistent with the 2D percolation exponent. The susceptibility data in Fig. 7 (b) are also found to collapse on a master-curve and show a crossover in the scaling exponent for the variation of $\chi_{\text{rig}}$ as a function of $|\phi-\phi_{c}|$. Far from the critical packing fraction ($|\phi-\phi_{c}|\gg 0$) $\chi_{\text{rig}}$ is found to scale with a power law exponent of $\gamma\approx 3.75$, as shown by the dotted line. Whereas, close to the critical point (i.e. $|\phi-\phi_{c}|<10^{-2}$) the scaling exponent ($\gamma$) is much lower, $\gamma\approx 0.3$, as shown in the main plot of Fig. 7 (b) for $N=1000$. However, it is noteworthy that for the larger system size, i.e. $N=8000$, the data show a crossover in the scaling exponent from $\gamma\approx 3.75$ to $\gamma\approx 2.4$, as presented in the inset of Fig. 7 (b). Notably, $\gamma=2.4$ is in agreement with the critical exponent for susceptibility, $\gamma=43/18\approx 2.4$, in 2D percolation transition [46, 63]. Santra et al. [56] have carried out a similar analysis to show that the scaling exponent $\gamma$ is approximately equal to $7/4$ close to the critical transition, consistent with the 2D Ising model. However, they have extracted such a value for a relatively smaller system size. From the present analysis it is clear that with increasing system size the susceptibility exponent for the rigidity transition approaches towards a value consistent with the critical exponent in 2D percolation. Furthermore, we have carried out finite size scaling analysis to show that the 2D percolation exponents for the order-parameter and susceptiblity actually work very well by demonstrating a good data collapse for the scaling functions, which is discussed in the subsequent section.
Nevertheless, the key observation from the results presented in Fig. 7 (a) and 7 (b) indicates that shear thickening and jamming behavior in non-Brownian suspensions remain unaffected by the polydispersity and shape of the particle size distribution. This has major implications on the understanding of the microscopic flow behavior of polydisperse suspensions since we could predict the shear thickening and shear jamming limits by appropriately rescaling the data.
@@ IV.2 Finite size scaling
Critical phase transitions in equilibrium systems are always associated with a diverging correlation length [52, 70, 10, 7] which results into the universal scaling relations for the order-parameter and susceptibility as discussed in the previous section. While our systems of shear thickening suspensions are fundamentally out of equilibrium, the observation of a critical transition for such a system is in itself a noteworthy finding. Furthermore, the agreement with the scaling exponents of the 2D percolation transition suggests that the fraction of particles in rigid clusters ($f_{\text{rig}}$) can be used as an order-parameter analogous to the percolation probability (which is the typical order-parameter for percolation transition) to explain the non-equilibrium phase transition close to the shear jamming in dense suspensions, where the system undergoes a transition from a flowable to a non-flowable state. While approaching a critical point, the correlation length ($\xi$), associated with the order parameter, is found to increase as $\xi\sim|t|^{-\nu}$, where $\nu$ is a critical exponent and $|t|$ is defined as the relative distance from the critical point in the limit of an infinite system size, i.e. in the context of the present study $|t|=|\phi-\phi_{c_{\infty}}|$, where $\phi_{c_{\infty}}$ is the critical packing fraction for an infinitely large system [52, 10]. This implies that for a system with linear dimension $L$, the correlation length at the critical transition becomes $\xi=L$, as $L\rightarrow\infty$. For finite size systems, the divergence of the correlation length near the critical point is characterized by the following finite size scaling relations [52, 70, 10, 4, 20, 67],
$\xi(t,L)$ $=$ $L\mathcal{G}(t\,L^{1/\nu})$ (15)
$f(t,L)$ $=$ $L^{-\beta/\nu}\mathcal{E}(t\,L^{1/\nu})$ (16)
$\chi(t,L)$ $=$ $L^{\gamma/\nu}\mathcal{F}(t\,L^{1/\nu})$ (17)
Here, $\mathcal{G}$, $\mathcal{E}$ and $\mathcal{F}$ are the scaling functions for the correlation length, order-parameter and susceptibility, respectively, and $t\,L^{1/\nu}$ is the corresponding scaling variable. This allows us to determine the critical exponents, $\beta$, $\gamma$ and $\nu$, from simulations of systems with finite domain and to identify the characteristics of the critical phenomena. Here, we have investigated finite size scaling behavior of polydisperse suspensions near the rigidity transition by simulating systems with different number of particles ($N=1000$, $2000$, $3000$ and $8000$) at stress $\sigma=100$. In our 2D simulations, since $N\propto L^{2}$, the system size is defined by $\sqrt{N}$. The correlation length for the rigid clusters can be estimated from a length scale associated with the cluster size. We have evaluated a length scale from the decay of the pair-correlation function, $g(r)$, for the rigid particles (identified by the pebble game algorithm) and used it as an estimate of the correlation length $\xi$ at different system sizes. This is illustrated in Figs. 8, where the variations of $g(r)$ are shown for $\phi=0.78$ and $\phi=0.794$, at different system sizes. It is to be noted that for any given value of $\phi$, the pair-correlation function is normalized by the corresponding number density of rigid particles only (i.e. non-rigid particles are excluded from the calculation). The long-range correlation length scale is extracted from the decay of the pair-correlation function by a method to identify the value of $r$ (radial distance) at which $g(r)$ settles to 1, which corresponds to the average number density of rigid particles. In this approach we have fitted the tail of the decay of g(r)1g(r)-1 by an exponential function of the form $y=A\exp(B\,r)$, and estimated the value of $r$ at which g(r)1g(r)-1 becomes $0.01$ (i.e. when the correlation function has decayed by 9999%) and considered that as an estimate of the correlation length ($\xi$). Uncertainty on the extracted value of $\xi$ is evaluated from the error estimate of the fitting parameters $A$ and $B$ by propagation of error. Insets in Fig. 8 display the variation of g(r)1g(r)-1 as a function of $r$ on a semi-log scale, which has been used to fit the exponential function as discussed above. While this approach works very well for the larger system size, close to the critical transition (i.e. $(\phi_{c_{\infty}}-\phi)<10^{-2}$, for $\phi<\phi_{c_{\infty}}$, where $\phi_{c_{\infty}}\approx 0.7946$ is evaluated by extrapolation of the $\phi_{c}$ values for $N\rightarrow\infty$), the exponential fit may produce large errorbars for the fitting parameters in case of smaller systems and far from $\phi_{c_{\infty}}$, where there are more fluctuations in $g(r)$. However, for the finite size scaling analysis it is necessary to estimate the correlation length close to the critical packing fraction, which could be estimated with negligible errorbars as presented in Fig. 9 (a). Notably, the correlation length ($\xi$) extracted in this manner is found to increase with increasing system size and packing fraction. This suggests that there is a divergence in the correlation length as the packing fraction approaches the critical point ($\phi_{c_{\infty}}$) from below and at $\phi=\phi_{c_{\infty}}$, the correlation length $\xi\rightarrow\infty$ in the limit of infinite system size.
@@
Figure 8: Pair-correlation function for rigid particles at (a) $\phi=0.78$ and (b) $0.794$ for normally distributed polydisperse systems ($\alpha=0.133$) with number of particles $N=2000$, $3000$ and $8000$. All the data are computed at stress $\sigma=100$. Insets show the variation of g(r)1g(r)-1 as a function of $r$ on a semi-log scale. The solid line in the inset represents an exponential fit of the form $g(r)-1=A\exp(B\,r)$.
@@
The finite size scaling relations in Eqs. (15)-(17) are tested for the normally-distributed polydisperse suspensions (with $\alpha=0.133$) by computing $\xi$, $f_{\text{rig}}$ and $\chi_{\text{rig}}$ for systems with N=1000, 2000, 3000N=1000,\,2000,\,3000 and $8000$ particles. Fig. 9 (a) presents the scaling of $\xi/L$ as a function of the scaling variable $|t|\,L^{1/\nu}$, where $t=(\phi-\phi_{c_{\infty}})$ and $\phi_{c_{\infty}}\approx 0.7946$. A good data collapse is observed for $\xi/L$ by choosing a value of the scaling exponent $\nu\approx 1.33$. It should be noted that the unscaled data are well separated, as presented in the inset of Fig. 9 (a), validating the choice of the scaling variables and parameters for obtaining the data collapse. As shown in Fig. 9 (b), the scaling of $\chi_{\text{rig}}/L^{\gamma/\nu}$ also indicates a data collapse for the different system sizes close to the critical transition. Notably, for $\chi_{\text{rig}}$ the quality of data collapse near the critical point (i.e. $|t|\rightarrow 0$) improves with increasing system size. The corresponding values of the scaling exponents are chosen to be $\gamma\approx 2.4$ and $\nu\approx 1.33$, which are consistent with the susceptibility scaling exponent found in the cross-over regime near the critical point for $N=8000$ particles, as shown in the inset of Fig. 7 (b), and finite size scaling of correlation length in Fig. 9 (a). It is noteworthy that $\nu=4/3\approx 1.33$ is in agreement with the value of the critical exponent in 2D percolation transition [63, 46]. Similarly, the finite size scaling for the order-parameter $f_{\text{rig}}$ is demonstrated in Fig. 9 (c), where a collapse of data is observed near the critical point with $\beta\approx 0.14$ and $\nu\approx 1.33$. It is to be noted that there are two branches of the scaling function $f_{\text{rig}}L^{\beta/\nu}$, where the upper and the lower branch represents the data for $\phi>\phi_{c_{\infty}}$ and $\phi<\phi_{c_{\infty}}$, respectively. The inset in Fig. 9 (b) and 9 (c) presents the corresponding unscaled data for $\chi_{\text{rig}}$ and $f_{\text{rig}}$, respectively. Moreover, similar to the scaling of the order-parameter and susceptibility discussed in section IV.1, the consistency in the values of exponent $\beta$, $\gamma$ and $\nu$ with the exponents of 2D percolation transition suggests that the rigidity transition in shear thickening suspensions can be correlated to a percolation transition [66, 22].
Refer to caption Refer to caption (a) (b) Refer to caption (c)